OpenStudy (anonymous):

The width and the length of a rectangle are consecutive even integers. If the width is decreased by 3 inches, then the area of the resulting rectangle is 24 square inches. What is the area of the original rectangle?

OpenStudy (zpupster):

area= l*W width = 2w lenghth=2w+2 consecutive even integer decrease width by 3 2w-3 area = (2w+2)(2w-3)=24 4w^2 - 6w + 4w -6 = 24 subtract 24 both sides 4w^2 -2W -30 = 0 FIND THE ZEROS hint use quadratic formula IGNORE ANY NEGATIVE NUMBERS plug back in original then compute area